OpenStudy (anonymous):

If sin (A) = -12/13 with A in QIII, find cos (A/2).

OpenStudy (anonymous):

5/13 in quadrants I and IIII and negative in the other two quadrants.

OpenStudy (anonymous):

Use cos(B/2) = ± sqrt([1 + cos B] / 2) after finding cos

OpenStudy (cwrw238):

cos (A) = -5/13

OpenStudy (cwrw238):

cos (A) = 2 cos^2(A/2) - 1 -5/13 = 2 cos^2(A/2) - 1 solve for cos( A/2)

OpenStudy (cwrw238):

2 cos^2(A/2 = 8/13 cos^2(A/2) = 4/13 cos (A/2) = +- 2 / sqrt 13 the value must be negative (cos is neg in Q3) so cos(A/2) = -2/sqrt13